A coffee company wants a new flavor of Cajun coffee. How many pounds of coffee worth $7 a pound should be added to 30 pounds of coffee worth $2 a pound to get a mixture worth $5 a pound
Answer by mananth(16946)

--- price ---------------- Amount
coffee I 2 ---------------- 30 lbs
coffee II 7 ---------------- x lbs
Mixture 5 ---------------- 30 + x lbs

2 * 30 + 7 x = 5 ( 30 + x )
60 + 7 x = 150 + 5 x
7 x -5 x = -60 + 150
2 x = 90
/ 2
x = 45 lbs coffee II

m.ananth@hotmail.ca